The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs. Every solution that I come up with is not a whole number or does not work when you try to check it.
Answer by Paul(988)

Since its issolces it ust be with 2 sides of the given formula = 4x-8
And the originial base which is x
4x-8+x=29
5x=37
x=7.4
4(7.4)-8=21.6
21.6/2=10.8

Hence, the base is 7.4cm and the length is 10.8 each.
Paul.